Sara Holliday must earn a return of 7 % on an investment that requires an initial outlay of $2,400 and promises to return $6,200 in 11 years.

a. Use present-value techniques to estimate the yield on this investment.

b. On the basis of your finding in part a , should Sara make the proposed investment? Explain.

a. The yield on this investment is
Answer #1

a) PV = FV/(1+r)11
Yield on this investment (r) = (FV/PV)1/11-1 = 9.01%

b)Yes Sara must make the proposed investment since yield of 9.01% is greater than 7%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well